a nozzle is supplied with steam having a specific enthalpy of 3268kj/kg, specific volume of 0.4m^3/kg and velocity of 10m/s. at exit of the nozzle, the specific volume and specific enthalpy of the steam are 1.32m^3/kg and 3072kj/kg, respectively while losing heat at rate of 25kw . for an inlet area of 800cm^2, determine : 1.the mass rate flow

2.the exit velocity

3.the change in kinetic energy

4.exit volume flow rate



1
Expert's answer
2021-06-08T09:13:03-0400

1.The mass rate flow

"\dot{m}=\rho Ac=\frac{1}{v}Ac= \frac{1}{1.32}*\frac{800}{10000}*10= 0.61 kg/s"

2.The exit velocity

"c_2=\sqrt{2(h_1-h_2)}=\sqrt{2(3268-3072)}=19.799 m/s"

3.The change in kinetic energy

"KE_2-KE_1 = \frac{1}{2}mv_2^2-\frac{1}{2}mv_1^2=\frac{1}{2}m(v_2^2-v_1^2)=0.5*0.61(19.799^2-10^2)"

"\triangle KE=0.5*0.61(19.799^2-10^2) =89.06 J"

4.The exit volume flow rate

"Q=vA=19.799*\frac{800}{10000}=1.58392 m^3/s"